Kurtulus Corporation uses the weighted-average method in its process costing system. Data concerning the first processing department for the most recent month are listed below:

Beginning work in process inventory:
Units in beginning work in process inventory 1,500
Materials costs $ 7,900
Conversion costs $ 3,200
Percent complete with respect to materials 55 %
Percent complete with respect to conversion 25 %
Units started into production during the month 7,400
Units transferred to the next department during the month 6,600
Materials costs added during the month $ 111,000
Conversion costs added during the month $ 84,100
Ending work in process inventory:
Units in ending work in process inventory 2,300
Percent complete with respect to materials 70 %
Percent complete with respect to conversion 55 %

The total cost transferred from the first processing department to the next processing department during the month is closest to:

Garrison 16e Rechecks 2017-08-28

Multiple Choice

  • $185,209

  • $168,841

  • $210,199

  • $194,509
Answer #1

Equivalent unit of material = 6600+(2300*70%) = 8210

Equivalent unit of conversion = 6600+(2300*55%) = 7865

Cost per equivalent unit of material = (7900+111000)/8210 = 14.482

Cost per equivalent unit of conversion = (3200+84100)/7865 = 11.100

Cost of unit transferred out = (14.482+11.10)*6600 = 168841

So answer is b) $168841